OpenStudy (anonymous):

a 50 gram sample of a substance that's used to treat thyroid disorders has a ka-value of 0.1133. Find the substances half-life, in days. Round your answer to the nearest tenth.

OpenStudy (irishboy123):

\(\large N(t) = N_o \ e^{-k \ t}\) \(\large N(t_{half}) = \frac{1}{2}N_o = N_o \ e^{-k \ t _{half}}\) \(\large ln(\frac{1}{2}) = - k \ \ t _{half} \) \(\large t_{half} = \frac{ln(\frac{1}{2})}{-k} = \frac{ln(2)}{k}\)